AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 20 of 33  ELOG logo
ID Date Author Subject Text Attachments
  279   Fri Feb 25 06:57:09 2022 George, EstherCurrent status08:00
Beam seems to be back
Intensity ~7.8e7
 2022-02-25_08-03-21leakage.png 
  278   Fri Feb 25 06:28:12 2022 George, EstherCurrent Status07:30
Some signals on the beam monitor,
seems they are optimising it.
 2022-02-25_07-33-44exRate.png2022-02-25_07-35-11xyTotal.png 
  277   Fri Feb 25 05:04:16 2022 Rui-Jiu Chen, AlexcheckStill  no beam in CRING.

Motor fully out position.

SW3
 14x 
  276   Fri Feb 25 03:20:29 2022 Rui-Jiu Chen, AlexCurrent statusGleb Vorobyev tried for few hours to fix
the ion source but without success.

So, there will be no beam in CRYRING
  
  275   Fri Feb 25 03:01:11 2022 Rui-Jiu Chen, AlexcheckStill  no beam in CRING.

Motor fully out position.

SW3
 14x 
  274   Fri Feb 25 01:12:49 2022 Rui-Jiu Chen, Alex Check Still  no beam in CRING.

Motor fully out position.

SW3
 13x 
  273   Fri Feb 25 00:00:40 2022 Rui-Jiu Chen, AlexCheckThere are some problems with ion source. 
No beam in CRYING.

Motor fully out position.
 14x 
  272   Thu Feb 24 23:15:25 2022 Rui-Jiu Chen, AlexProblems with ion souce.    
  271   Thu Feb 24 22:45:08 2022 Jordan, George, Rui-Jiu Chencheck- Motor fully out position.

SW3

- Merger OK
 14x 
  270   Thu Feb 24 14:31:25 2022 JM, ML, CNBeam Tuning Run 7615:30

1.5 MeV/u beam currently being
tuned
 6x 
  269   Thu Feb 24 09:43:16 2022 JG, MScheckBeam is still beeing tuned. Target is off.

 

- Motor on manual
 13x 
  268   Thu Feb 24 08:30:36 2022 JG, MSAIDA5 calibration test failed - solvedUPDATE: THE PROBLEM IS SOLVED FOR NOW:

Doing the calib for the 5th time
eventually worked.
 Screenshot_from_2022-02-24_09-09-36.pngScreenshot_from_2022-02-24_09-29-24.png 
  267   Thu Feb 24 08:15:23 2022 JG, MScheckBeam is still beeing tuned. Target is off.

There is a problem with ADC calibration
of aida5: see entry https://elog.ph.ed.ac.uk/CARME/268
 13x 
  266   Thu Feb 24 07:25:53 2022 OH (Remote)Recover of system after crash08:00 Woke up to message from Carlo explaining
the server had crashed.
      Anydesked into the machine.
  
  265   Thu Feb 24 05:09:47 2022 CB, PE, EMInlet valve at target closed + problems with turbopumpThe turbo pump E1 is continuing to draw
too much current and the rotation frequency
has steadily decreased from 817 Hz at the
  
  264   Thu Feb 24 04:17:35 2022 CB, TD (remotely)MIDAS DAQ no longer workingDAQ server crashed around 24:00

Rebooted server but could not could
not rescue remotely - could not contact raspberry
 Bad_Time_2022-02-24_04-34-12.pngBadTime_WRTime_2022-02-24_04-34-53.png 
  263   Wed Feb 23 22:29:50 2022 PE, EM, CB (remotely)check - 23:30Motor position ok

Data rate: 1030kbytes/s

T ok
  
  262   Wed Feb 23 21:27:29 2022 TD, CNcheck Motor position ok

Data rate: 1030kbytes/s

T ok
 11x 
  261   Wed Feb 23 20:52:56 2022 CN, TDSystem wide checks
Attachments 1-5 aida01-aida05 options
 11x 
  260   Wed Feb 23 20:06:43 2022 TD CNE141 kinematics
Assuming DSSSD is 1000um thick we can expect
to observe protons for E_lab=11MeV as follows
 E141_kinematics.txt 
ELOG V3.1.4-unknown